OpenStudy (anonymous):

HELP :( Chelsea has $45 to spend at the fair. She spends $20 on admission and $15 on smacks. She wants to play a game that costs $0.65 per game. Write an inequality to find the maximum number of times, x, Chelsea can play the game? a) Using this inequality, determine the maximum number of times she can play the game.

OpenStudy (anonymous):

\[45\ge 35+.65x\]

OpenStudy (anonymous):

So to find the maximum # of times she can play... Subtract 35 on both sides. \[10\ge .65x\] Divide by .65 on both sides. \[15.384\ge x\] Then round down. She can play the game 15 times.

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Np.